# Direct Sampling from Conditional Distributions by Sequential Maximum Likelihood Estimations

### v2

@@ 1 Introduction
Consider a discrete sample with state space $[m]:=\{1,2,\ldots,m\}$ for a positive integer $m$. In this paper, we will discuss the discrete exponential families in statistics, which are also called log-affine models or toric models.
@@
Let $A=(a_{ij})$ $\in\mathbb{Z}^{d\times m}$ be a matrix of integers such that no row or column is zero vector and $(1,\ldots,1)\in{\rm rowspan}(A)$. Let $x\in\mathbb{R}^{m}_{>0}$. The log-affine model associated with the configuration matrix $A$ is the set of probability distributions
@@
In this paper, we consider the closure, called the extended log-affine model by [11], Section 4.2.3, whose support may be at the boundary of the simplex. This extension ensures the existence of the maximum likelihood estimate (MLE), whether or not observed counts satisfy constraints to be positive.
@@
Since the discrete exponential family is the standard family of discrete distributions, sampling from the conditional distribution (3) has various statistical demands. However, it also has been considered usually impossible (see, e.g. [20], Chapter 9). This motivates us to use the Metropolis algorithm, originally devised for sampling from canonical distributions in physics and one of the most common tools in the Markov chain Monte Carlo methods. In the Metropolis algorithm, the unique stationary distribution of an ergodic Markov chain is the distribution we need (see, e.g., [13], Chapter 3). We call the chain Metropolis chain. Diaconis and Sturmfels [6] proposed by using Gröbner bases of the toric ideal of a configuration matrix to construct a basis of moves in the Metropolis chain. They called such a basis, or a generating set for the toric ideal, a Markov Basis. The state space of the chain is the $b$-fiber (4), and the ratio of two monomials, which correspond to the current state and the proposed next state, is computed in each step of the chain.
@@
The direct sampling algorithm is a random walk or a Markov chain on a bounded integer lattice defined below. One sample path of the Markov chain is equivalent to one contingency table. Note that while the sequential draw of the contingency table is the same as in Chen et al. [2], as Markov chains the two are quite different and the units of the direct sampling are individual counts rather than cells. Figure 1 is the Hasse diagram of an example.
Figure 1: The Markov lattice for the matrix in Example 2.8 with the maximum $b={}^{t}(1,2,2,1)$. A sample path is shown by solid edges.
@@
and the maximum and the minimum are $b$ and 00, respectively. Here, $a_{j}$ denotes the $j$-th column vector of $A$.
@@
More generally, the main obstacle common to using the Markov basis and the direct sampling algorithm for general models is the computational cost of Gröbner bases. The Gröbner basis guarantees the realizability of these methods, however, Buchberger’s algorithm for computing Gröbner bases is impractical since it demands a huge computational cost (see, e.g., [4], Section 2.9). Therefore, in practice, we have to avoid using Gröbner bases. The challenge is to obtain Markov bases or connection matrices by some consideration using properties of each model, rather than resorting to Buchberger’s algorithm. A book-length treatment of this issue for the Markov bases is [1] and for the direct sampling algorithm is [15].
@@
In line with this trend, special MLEs have been attracting attention in recent years. Discrete statistical models for which the MLE admits an expression of a rational function of the observed counts are called rational, and they were characterized by [10] and [7]. Such an expression of the MLE is called the rational MLE. The MLE of the expected count of decomposable graphical models is given by a rational function of the observed counts (see, e.g., [11], Section 4.4), and it was shown that a log-linear model is rational if and only if it is a decomposable graphical model [8]. In this paper, we will see that the UMVUE and the MLE coincide if and only if a log-linear model is a decomposable graphical model.
We restate the observation above about the direct sampling algorithm as follows: if a model is log-linear and rational, then the computation of the transition probability (6) is immediate, and the direct sampling algorithm works efficiently and exactly. Otherwise, we may replace the UMVUE with the MLE and run the algorithm using existing methods to compute the MLE, such as the iterative proportional scaling algorithm, with the expense of bias of the MLE, that asymptotically disappears because of its consistency.
This paper is structured as follows. In Section 2, we introduce the details of the direct sampling algorithm proposed in [14] and show that the transition probability is the ratio of the UMVUE of the expected count to the total number of counts. Then, after summarizing the relevant properties of the MLE, we show that the UMVUE and the MLE coincide if and only if a log-linear model is a decomposable graphical model and the resulting MLE is rational. Lastly, we discuss the use of the MLE and reveal the key how we avoid computation of the connection matrices of the holonomic ideal. Section 3 is devoted to numerical experiments. Using several contingency tables as examples, we demonstrate how the direct sampling algorithm works. Section 4 summarizes the significance of this paper.
To the best of author’s knowledge, this is the first paper to demonstrate direct sampling from the conditional distribution of the discrete exponential family in a feasible time, although it is generally approximate in the sense that it is only asymptotically exact, not exact.
@@ 2 Algorithm
  • Step 1: Set $t=1$, $\beta=b$, and $\nu=n$.
  • Step 2: Compute $P(\beta,\beta-a_{j};x)$ for each $j\in[m]$ in (6).
@@
  • Step 4: If $\nu=1$, go to Step 5. Else, set $t\leftarrow t+1$, $\beta\leftarrow\beta-a_{j_{t}}$, $\nu\leftarrow\nu-1$, and go to Step 2.
@@
A merit of Algorithm 1 is that we do not have to construct the Markov lattice $\mathcal{L}_{A}(b)$. A Markov transition occurs only between adjacent elements because if otherwise, $Z_{A}(\beta-a_{j};x)=0$ since $\beta-a_{j}\not\in\mathbb{N}A$. For the same reason, a sample path never hits outside of $\mathcal{L}_{A}(b)$, or more precisely, the set $({\rm Cone}(A)\cap\mathbb{Z}^{m})\setminus\mathbb{N}A$, where ${\rm Cone}(A):=\{Av:v\in\mathbb{R}^{m}_{\geq 0}\}$ is the polyhedral cone generated by the column vectors of $A$.
@@ 2.1 The MLE
$\sum_{j\in[m]}a_{ij}\mu_{j}:=\sum_{j\in[m]}a_{ij}\mathbb{E}[U_{j}]=\mathbb{E}[B_{i}]=\frac{\partial\psi}{\partial\xi_{i}},\quad i\in[d],$
where the last equality holds by $\mathbb{E}[\partial l/\partial\xi_{i}]=0$. The MLEs of the canonical parameters $\hat{\xi}_{i}$, $i\in[d]$ constitute the unique solution of the estimating equations $b_{i}=\partial\psi/\partial\xi_{i}$. Therefore, we have
@@
More precisely, we have
Theorem 2.2 ([11], Theorem 4.8).
@@
for a sample of the vector of sufficient statistics $b$ is unique. This holds under multinomial as well as Poisson sampling.
In what follows, we sometimes denote the vector of the MLE $\hat{\mu}$ by $\hat{\mu}(b;x)$ to explicitly show the dependence on sufficient statistics $b$ and $x$. We apply this rule to the expected counts $\mu$ and the UMVUE $\tilde{\mu}$ discussed below.
@@
  • Step 3’: Pick $j_{t}=j$ with probability $\hat{\mu}_{j}(\beta;x)/\nu$, $j\in[m]$,
@@
2.2 The UMVUE and Markov transitions
We recall some basic properties of the $A$-hypergeometric polynomial (5).
Definition 2.4.
$(\sum_{j\in[m]}x_{j}\partial_{j}-\deg(\beta))Z_{A}(\beta;x)=0,$ (10)
$\partial_{j}Z_{A}(\beta;x)=Z_{A}(\beta-a_{j};x),\quad j\in[m].$ (11)
These properties give
Lemma 2.5.
Proof.
By the conditional distribution (3), we observe that
$\mathbb{E}(U_{j}|AU=b)=\sum_{v\in\mathcal{F}_{b}(A)}v_{j}\frac{x^{v}}{v!}=\frac{\partial_{j}Z_{A}(b;x)}{Z_{A}(b;x)}x_{j}=\frac{Z_{A}(b-a_{j};x)}{Z_{A}(b;x)}x_{j},\quad j\in[m],$
where the last equality follows by the contiguity relation (11). This is the unbiased estimator of $\mu_{j}$ because $\mathbb{E}\{\mathbb{E}(U_{j}|AU=b)\}=\mathbb{E}[U_{j}]=\mu_{j}$. Since $b$ is the vector of complete sufficient statistics, by the Lehmann-Scheffé theorem (Theorem 2.1.11, [12]), it is the unique unbiased estimator that is a function of $b$. ∎
By the homogeneity (10), we have
Before closing this subsection, we mention an observation about the UMVUE. Note that the following proposition does not mean that the UMVUE always coincides with the MLE since the UMVUE generally does not lie in the model manifold $\mathcal{M}_{A}$. If it lies, it should coincide with the MLE by Theorem 2.2. We will be back to this issue in Section 2.3.
@@ 2.3 The UMVUE and the rational MLE
Here, we present a small example in [10, 7] in our context.
Example 2.8 (Two-way contingency tables of the independence model).
respectively. The conditional distribution is
@@
where the rows are for $u_{1\cdot}$, $u_{2\cdot}$, $|u|$, $u_{\cdot 1}$, and $u_{\cdot 2}$, respectively. The Markov lattice $\mathcal{L}_{A}(b)$ with $b={}^{t}(1,2,2,1)$ is in Figure 1, and the transition probability of Step 3’ of Algorithm 1 becomes
$\frac{\hat{\mu}(b;1)}{|u|}=\frac{u_{i\cdot}}{|u|}\frac{u_{\cdot j}}{|u|},\quad i,j\in[2],$ (14)
where $a_{ij}$ denotes the column vector of $A$ cording the $(i,j)$-cell of the table.
In the remainder of this subsection we concentrate on the log-lineal model, so we set $x=1$.
Geiger et al. [8] showed that a log-linear model has the rational MLE if and only if it is a decomposable graphical model (see [11] for the definition of the decomposable graphical model). Using this fact, we have the following theorem.
The notation for graphical models can be complex; we follow that of [11] here. The correspondence between this notation and the notation of the $A$-hypergeometric polynomial is explained in detail in Section 3.2 of [16], so we will not repeat it here. The two-way contingency table of the independence model discussed above is an example of a decomposable graphical model: the undirected graph consists of two vertices corresponding to the row and column of the contingency table and no edge. The two vertices are the cliques, and the separator is empty.
@@
The UMVUE and the MLE of a log-linear model coincides if and only if the model is a decomposable graphical model. Moreover, the MLE is rational.
@@
where $\mathcal{S}$ and $\mathcal{C}$ are the sets of separators and cliques of the undirected graph $G$, respectively, and $\nu(S)$ is the number of appearances of separator $S$ in a perfect sequence of the cliques. The index $j$ specifies a state of the vertex set of $G$. The indices $j_{S}$ and $j_{C}$ are indices obtained by summing up indices of $j$ other than $S$ and $C$, respectively, and $u(j_{S})$ and $u(j_{C})$ are the number of counts of states consistent with $j_{S}$ and $j_{C}$, respectively. Lemma 2.5 shows that the UMVUE is
@@
$=\frac{\prod_{C\in\mathcal{C}}u(j_{C})}{\prod_{S\in\mathcal{S}}u(j_{S})^{\nu(S)}}.$
@@
(only if) Suppose we have a nondecomposable graphical model. Then, Theorem 4.4 of [8] implies that we cannot always find a rational solution to the maximum likelihood estimation. On the other hand, the expression of the UMVUE in Lemma 2.5 implies that the UMVUE is always a rational number, because of the definition of the $A$-hypergeometric polynomial (5). Therefore, the MLE and the UMVUE never coincide. ∎
With Proposition 2.6, we immediately have
@@
The UMVUE of a log-linear model lies in the model manifold $\mathcal{M}_{A}$ if and only if the model is a decomposable graphical model.

### v3

@@ 1 Introduction
Consider a discrete sample with state space $m]:=\{1,2,\ldots,m\}$ for a positive integer $m$. In this paper, we will discuss the discrete exponential families in statistics, which are also called log-affine models or toric models.
@@
Let $A=(a_{ij})$ $\in\mathbb{Z}^{d\times m}$ be a matrix of integers such that no row or column is the zero vector and $(1,\ldots,1)\in{\rm rowspan}(A)$. Let $x\in\mathbb{R}^{m}_{>0}$. The log-affine model associated with the configuration matrix $A$ is the set of probability distributions
@@
In this paper, we consider the closure, called the extended log-affine model by [12, Section 4.2.3], whose support may be at the boundary of the simplex. This extension ensures the existence of the maximum likelihood estimate (MLE), whether or not observed counts satisfy constraints to be positive.
@@
Since the discrete exponential family is the standard family of discrete distributions, sampling from the conditional distribution (3) has various statistical demands. A typical use in statistics is hypothesis testing. To evaluate the significance of our data under a hypothetical model, we generate random variables called statistics from its distribution and estimate the significance by the proportion of generated values greater than the observed value. When evaluating power of a test, we generate random variables from a distribution under a certain hypothetical model and estimate the power by the proportion of generated values that exceed a given significance level. Such hypothetical models include the non-independence model of two-way contingency tables and the no-three-way interaction model for the three-way contingency tables. However, sampling from such models has also been considered to be usually impossible. (see, e.g. [22, Chapter 9]). This motivates us to use the Metropolis algorithm, originally devised for sampling from canonical distributions in physics and one of the most common tools in the Markov chain Monte Carlo methods. In the Metropolis algorithm, the unique stationary distribution of an ergodic Markov chain is the distribution we need (see, e.g., [14, Chapter 3]). We call the chain Metropolis chain. Diaconis and Sturmfels [7] proposed to construct a basis of moves in the Metropolis chain by using Gröbner bases of the toric ideal of a configuration matrix. They called such a basis, or a generating set for the toric ideal, a Markov basis. The state space of the chain is the $b$-fiber (4), and the ratio of two monomials, which correspond to the current state and the proposed next state, is computed in each step of the chain.
@@
The direct sampling algorithm is a random walk or a Markov chain on a bounded integer lattice defined below. One sample path of the Markov chain is equivalent to one contingency table. See Example 1.3. Note that while the sequential draw of the contingency table is the same as in Chen et al. [2], as Markov chains the two are quite different and the units of the direct sampling are individual counts rather than cells.
@@
and the maximum and the minimum are $b$ and 0, respectively. Here, $a_{j}$ denotes the $j$-th column vector of $A$.
Example 1.3 (Two-way contingency tables of the independence model).
respectively. The conditional distribution is
Figure 1 is the Hasse diagram of an example.
Figure 1: The Markov lattice for the matrix in Example 1.3 with the maximum $b=(1,2,2,1)^{\top}$. A sample path is shown by solid edges.
Transitions occur along the edges. The sample path, shown with solid edges, is equal to one contingency table with sufficient statistics $b=(1,2,2,1)^{\top}$. In the first line of the following tables, asterisks mark the cells picked during each solid edge; each time a cell is picked, the corresponding marginal count drops by one. The second line shows how the contingency table is created by listing these picked counts. The general form of this process is given in Algorithm 1.
$\begin{array}[]{cc|c}*&&1\\ &&2\\ \hline\cr 2&1\end{array}$ $\quad\to$ $\begin{array}[]{cc|c}&&0\\ &*&2\\ \hline\cr 1&1\end{array}$ $\quad\to$ $\begin{array}[]{cc|c}&&0\\ &&1\\ \hline\cr 1&0\end{array}$ $\quad\to$ $\begin{array}[]{cc|c}&&0\\ &&0\\ \hline\cr 0&0\end{array}$ (18)
[…]
@@
More generally, the main obstacle common to using the Markov basis and the direct sampling algorithm for general models is the computational cost of Gröbner bases. The Gröbner basis guarantees the existence of a Markov basis, however, Buchberger’s algorithm for computing Gröbner bases is impractical since it demands a huge computational cost (see, e.g., [5, Section 2.9]). Therefore, in practice, we have to avoid using Gröbner bases. The challenge is to obtain Markov bases or connection matrices by some consideration using properties of each model, rather than resorting to Buchberger’s algorithm. A book-length treatment of this issue for the Markov bases is [1] and for the direct sampling algorithm is [16].
@@
In line with this trend, special MLEs have been attracting attention in recent years. Discrete statistical models for which the MLE admits an expression of a rational function of the observed counts are called rational, and they were characterized by [11] and [8]. Such an expression of the MLE is called the rational MLE. The MLE of the expected count of decomposable graphical models is given by a rational function of the observed counts (see, e.g., [12, Section 4.4]), and it was shown that an undirected discrete graphical model is rational if and only if it is a decomposable graphical model [9]. In this paper, we will see that the UMVUE and the MLE of an undirected discrete graphical model coincide if and only if it is a decomposable graphical model.
We restate the observation above about the direct sampling algorithm as follows: if a model is log-linear and rational, then the computation of the transition probability (29) is immediate, and the direct sampling algorithm works efficiently and exactly. Otherwise, we may replace the UMVUE with the MLE and run the algorithm using existing methods to compute the MLE, such as the iterative proportional scaling algorithm, with the expense of bias of the MLE, that asymptotically disappears (See Remark 2.16).
This paper is structured as follows. In Section 2, we introduce the details of the direct sampling algorithm proposed in [15] and show that the transition probability is the ratio of the UMVUE of the expected count to the total number of counts. Then, after summarizing the relevant properties of the MLE, we show that the UMVUE and the MLE of an undirected discrete graphical model coincide if and only if is a decomposable graphical model and the resulting MLE is rational. Lastly, we discuss the use of the MLE and reveal the key how we avoid computation of the connection matrices of the holonomic ideal. Section 3 is devoted to numerical experiments. Using several contingency tables as examples, we demonstrate how the direct sampling algorithm works. Section 4 summarizes the significance of this paper.
To the best of author’s knowledge, this is the first paper to demonstrate direct sampling from the conditional distribution of the discrete exponential family in a feasible time, although it is generally approximate.
@@ 2 Algorithm
  • Step 1: Set $\beta=b$.
    For $t=1,2,\ldots,n$ do
  • Step 2: Compute $P(\beta,\beta-a_{j};x)$ in (29) for each $j\in[m]$.
@@
  • Step 4: Set $\beta\leftarrow\beta-a_{j_{t}}$.
    End for
@@
A merit of Algorithm 1 is that we do not have to construct the Markov lattice $\mathcal{L}_{A}(b)$. A Markov transition occurs only between adjacent elements because otherwise, $Z_{A}(\beta-a_{j};x)=0$ since $\beta-a_{j}\not\in\mathbb{N}A$. For the same reason, a sample path never hits outside of $\mathcal{L}_{A}(b)$, or more precisely, the set $({\rm Cone}(A)\cap\mathbb{Z}^{m})\setminus\mathbb{N}A$, where ${\rm Cone}(A):=\{Av:v\in\mathbb{R}^{m}_{\geq 0}\}$ is the polyhedral cone generated by the column vectors of $A$.
2.1 Markov transitions and the UMVUE
We recall some basic properties of the $A$-hypergeometric polynomial (5). A book-length treatment of the $A$-hypergeometric functions is [21].
Definition 2.2.
$(\sum_{j\in[m]}x_{j}\partial_{j}-\deg(\beta))Z_{A}(\beta;x)=0,$ (32)
$\partial_{j}Z_{A}(\beta;x)=Z_{A}(\beta-a_{j};x),\quad j\in[m].$ (33)
These properties give the following lemma. In what follows, we sometimes denote the vector of the UMVUE $\tilde{\mu}$ by $\tilde{\mu}(b;x)$ to explicitly show the dependence on sufficient statistics $b$ and $x$. We apply this rule to the expected counts $\mu$ and the MLE $\hat{\mu}$ discussed below.
Lemma 2.3.
Proof.
By the conditional distribution (3), we observe that
$\mathbf{E}(U_{j}|AU=b)=\frac{1}{Z_{A}(b;x)}\sum_{v\in\mathcal{F}_{A}(b)}v_{j}\frac{x^{v}}{v!}=\frac{\partial_{j}Z_{A}(b;x)}{Z_{A}(b;x)}x_{j}=\frac{Z_{A}(b-a_{j};x)}{Z_{A}(b;x)}x_{j},\quad j\in[m],$
where the last equality follows by the contiguity relation (33). This is the unbiased estimator of $\mu_{j}$ because $\mathbf{E}[\mathbf{E}(U_{j}|AU=b)]=\mathbf{E}[U_{j}]=\mu_{j}$. Since $b$ is the vector of complete sufficient statistics, by the Lehmann-Scheffé theorem [13, Theorem 2.1.11], it is the unique unbiased estimator that is a function of $b$. ∎
By the homogeneity (32), we have
@@ 2.2 The MLE
$\sum_{j\in[m]}a_{ij}\mu_{j}:=\sum_{j\in[m]}a_{ij}\mathbf{E}[U_{j}]=\mathbf{E}[B_{i}]=\frac{\partial\psi}{\partial\xi_{i}},\quad i\in[d],$
where the last equality holds by $\mathbf{E}[\partial l/\partial\xi_{i}]=0$. The MLEs of the canonical parameters $\hat{\xi}_{i}$, $i\in[d]$ constitute the unique solution of the estimating equations
$\frac{\partial l(\xi;b)}{\partial\xi_{i}}=0,\quad\text{or}\quad b_{i}=\frac{\partial\psi}{\partial\xi_{i}},\quad i\in[d].$
Therefore, we have
@@
More precisely, we have the following theorem.
Theorem 2.4 ([12, Theorem 4.8]).
@@
for a sample of the vector of sufficient statistics $b$. Moreover, such MLE is unique. This holds under multinomial as well as Poisson sampling.
@@
  • Step 3’: Pick $j_{t}=j$ with probability $\hat{\mu}_{j}(\beta;x)/(n-t+1)$, $j\in[m]$,
@@
Before closing this subsection, we mention an observation about the UMVUE. Note that the following proposition does not mean that the UMVUE always coincides with the MLE since the UMVUE generally does not lie in the model manifold $\mathcal{M}_{A}$. If it lies in $\mathcal{M}_{A}$, it should coincide with the MLE by Theorem 2.4. We will be back to this issue in Section 2.3.
@@ 2.3 The UMVUE and the rational MLE
Here, we continue Example 1.3, which was discussed in [11, 8].
Example 2.8 (Two-way contingency tables of the independence model, cont.).
@@
where the rows are for $u_{1\cdot}$, $u_{2\cdot}$, $|u|$, $u_{\cdot 1}$, and $u_{\cdot 2}$, respectively. The transition probability of Step 3’ for the cell at the $j_{1}$-th row and $j_{2}$-th column (the $\{2(j_{1}-1)+j_{2}\}$-th column of configuration matrix $A$) becomes
$\frac{\hat{\mu}_{j_{1}j_{2}}(b;1)}{|u|}=\frac{u_{j_{1}\cdot}}{|u|}\frac{u_{\cdot j_{2}}}{|u|},\quad j_{1},j_{2}\in[2].$ (37)
This condition states that the manifold $\mathcal{M}_{A}$ (in the complex space) can be obtained as the Horn uniformization of an irreducible polynomial called the $A$-discriminant. Here, there exists a matrix $B\in\mathbb{Z}^{r\times l}$ of rank $r$, such that $BH=0$, which defines the $A$-discriminant. In this example,
$B=\begin{pmatrix}1&1&1&1&1\\ 0&0&1&2&2\end{pmatrix}$.
The matrix $B$ defines the set of polynomials in $t$: $F(t)=(x_{1}+x_{2})+x_{3}t+(x_{4}+x_{5})t^{2}$. The $A$-discriminant is an irreducible polynomial that defines the set of polynomials with multiple roots, $x_{3}^{2}-4(x_{1}+x_{2})(x_{4}+x_{5})$. See [11] for details. In the literature on the $A$-discriminant, we write this $B$ as $A$, but here we avoid a notational conflict. Note that the $A$ in the $A$-discriminant (here, it is denoted by $B$), and the $A$ in the $A$-hypergeometric system (Definition 2.2) are different.
In the remainder of this subsection we concentrate on the log-linear model, so we set $x=1$ in Definition 1.1.
Geiger et al. [9] showed that an undirected graphical model for discrete variables has a rational MLE if and only if it is a decomposable graphical model (an undirected graphical model for an undirected graph $G$ is the log-linear model in which the sufficient statistics correspond to the cliques of $G$. see [12] for the definition of the decomposable graphical model). Using this fact, we have the following theorem.
The notation for graphical models can be complex; we follow that of [12] here. The correspondence between this notation and the notation of the $A$-hypergeometric polynomial is explained in detail in [17, Section 3.2], so we will not repeat it here.
@@
The UMVUE and the MLE of an undirected discrete graphical model coincide if and only if the model is a decomposable graphical model. Moreover, the MLE is rational.
@@
where $\mathcal{S}$ and $\mathcal{C}$ are the sets of separators and cliques of the undirected graph $G$, respectively, and $\nu(S)$ is the number of appearances of separator $S$ in a perfect sequence of the cliques, and $b$ is the vector consisting of the sequence of $u(j_{C})$, $C\in\mathcal{C}$. The index $j$ specifies a state of the vertex set of $G$. The indices $j_{S}$ and $j_{C}$ are indices obtained by summing up indices of $j$ other than $S$ and $C$, respectively, and $u(j_{S})$ and $u(j_{C})$ are the number of counts of states consistent with $j_{S}$ and $j_{C}$, respectively. Lemma 2.3 shows that the UMVUE is
@@
$=\frac{\prod_{C\in\mathcal{C}}u(j_{C})}{\prod_{S\in\mathcal{S}}u(j_{S})^{\nu(S)}}.$ (38)
@@
(only if) Suppose we have a non-decomposable graphical model. Then, Theorem 4.4 of [9] implies that we cannot always find a rational solution to the maximum likelihood estimation. On the other hand, the expression of the UMVUE in Lemma 2.3 implies that the UMVUE is always a rational number, because of the definition of the $A$-hypergeometric polynomial (5). Therefore, the MLE and the UMVUE never coincide. ∎
With Proposition 2.6, we immediately have the following corollary.
@@
The UMVUE of an undirected discrete graphical model lies in the model manifold $\mathcal{M}_{A}$ if and only if the model is a decomposable graphical model.
Example 2.11 (Two-way contingency tables of the independence model, cont.).
The two-way contingency table of the independence model is a decomposable graphical model: the undirected graph consists of two vertices, say 1 and 2, corresponding to the row and column, respectively. It has no edges. A perfect sequence of the cliques is $C_{1}=\{1\}$, $C_{2}=\{2\}$ with separator $S=\emptyset$, $\nu(S)=1$. If we write $j=j_{1}j_{2}$, $j_{C_{1}}=j_{1}\cdot$ and $j_{C_{2}}=\cdot j_{2}$, we have $u(j_{C_{1}})=u_{j_{1}\cdot}$, $u(j_{C_{2}})=u_{\cdot j_{2}}$, and $u(S)=|u|$ (all indices are summed up). $b=(u_{1\cdot},u_{2\cdot},u_{\cdot 1},u_{\cdot 2})^{\top}$. Then, (38) gives the UMVUE
$\tilde{\mu}_{j_{1}j_{2}}(b;1)=\frac{u(j_{C_{1}})u(j_{C_{2}})}{u(S)}=\frac{u_{j_{1}\cdot}u_{\cdot j_{2}}}{|u|},$
which coincides with the rational MLE appeared in (37).
Note that Theorem 2.9 is a statement limited to undirected discrete graphical models, and does not provide a necessary and sufficient condition for the UMVUE and the MLE to coincide in log-linear models. In fact, as the following example shows, the UMVUE and the MLE can coincide in log-linear models that are not graphical models.
Example 2.12 (Two-way contingency tables of the quasi-independence models.).
Coons and Sullivant [4] classified the two-way contingency tables of the quasi-independence models (independence models with structural zeros) that have rational MLEs. The quasi-independence models are not graphical models, since the correspondence between the sufficient statistics and the cliques is incomplete. In Example 2.5 of [4], the rational MLE of the $3\times 3$ table whose cell at the third row and the third column is zero was given. The table is
$\begin{array}[]{ccc|c}u_{11}&u_{12}&u_{13}&u_{1\cdot}\\ u_{21}&u_{22}&u_{23}&u_{2\cdot}\\ u_{31}&u_{32}&0&u_{31}+u_{32}\\ \hline\cr u_{\cdot 1}&u_{\cdot 2}&u_{13}+u_{23}&|u|\end{array}.$
The configuration matrix, the vectors of the observed counts, and the vector of the sufficient statistics are
$A\!=\!\left(\begin{array}[]{cccccccc}1&1&1&0&0&0&0&0\\ 0&0&0&1&1&1&0&0\\ 0&0&0&0&0&0&1&1\\ 1&0&0&1&0&0&1&0\\ 0&1&0&0&1&0&0&1\\ 0&0&1&0&0&1&0&0\\ \end{array}\right),\quad u\!=\!\left(\begin{array}[]{c}u_{11}\\ u_{12}\\ u_{13}\\ u_{21}\\ u_{22}\\ u_{23}\\ u_{31}\\ u_{32}\end{array}\right),\quad{\rm and}\quad b=\left(\begin{array}[]{c}u_{1\cdot}\\ u_{2\cdot}\\ u_{31}+u_{32}\\ u_{\cdot 1}\\ u_{\cdot 2}\\ u_{13}+u_{23}\end{array}\right),$
respectively. The rational MLE of the expected count of the first row and the third column is
$\hat{\mu}_{13}(b;1)=\frac{u_{1\cdot}(u_{13}+u_{23})}{u_{1\cdot}+u_{2\cdot}},$
which coincides with the UMVUE. See Appendix for the details.